An engineer is going to redesign an ejection seat for an airplane. The seat was designed for pilots weighing between 130 lb and 181 lb. The new population of pilots has normally distributed weights with a mean of 140 lb and a standard deviation of 25.6 lb.
Answer #1

Solution :

Given that ,

mean = = 140

standard deviation = = 25.6

Using z table,  

= P(130< x < 181) = P[(130-140) / 25.6< (x -) / < (181-140) /25.6 )]

= P(-0.39 < Z < 1.60)

= P(Z < 1.60) - P(Z < -0.39)

Using z table,  

= 0.9452-0.3483

= 0.5969